# Consistency Models for Scalable and Fast Simulation-Based Inference

### v2

@@ 3 Consistency Model Posterior Estimation
Diffusion models have one crucial drawback: At inference time, they require solving many differential equations which slows down their sampling speed. This is particularly troublesome in SBI applications, which typically require thousands of samples for thousands of data sets. Consistency models (Song et al., 2023) address this problem with a new type of generative models that support both single-step and multi-step sampling. In the following, we extend the formulation of consistency models by Song et al. (2023) to accommodate conditioning information, such as data $\mathbf{x}$ or a jointly learned embedding $h(\mathbf{x})$ (Radev et al., 2020).
@@ 3.1 Conditional Consistency Models
The consistency function $f:(\boldsymbol{\theta}_{t},t;\,\mathbf{x})\mapsto\boldsymbol{\theta}_{\varepsilon}$ maps points across the solution trajectory $\{\boldsymbol{\theta}_{t}\}_{t\in[T,\varepsilon]}$ to the trajectory’s origin $\boldsymbol{\theta}_{\varepsilon}$ given a fixed conditioning variable $\mathbf{x}$ and the probability flow ODE in Eq. 4. To achieve this with established score-based diffusion model architectures, we can use a free-form neural network $F_{\boldsymbol{\phi}}(\boldsymbol{\theta},t;\,\mathbf{x})$ which is parameterized through skip connections of the form
@@
where $c_{\text{skip}}(t)$ and $c_{\text{out}}(t)$ are differentiable and fulfill the boundary conditions $c_{\text{skip}}(\varepsilon)=1$ and $c_{\text{out}}(\varepsilon)=0$. Consistency models are originally motivated as a distillation technique for diffusion models. However, Song et al. (2023) show that training consistency models in isolation is possible, and we base our method on their direct approach.
Once the consistency model has been trained, generating draws from the approximate posterior is straightforward by drawing samples from the noise distribution, $\boldsymbol{\theta}_{T}\sim\mathcal{N}(\mathbf{0},T^{2}\mathbf{I})$, which shall then be transformed into samples from the target distribution, like in a standard diffusion model. In contrast to diffusion models, however, we do not need to solve a sequence of differential equations for this transformation. Instead, we can use the learned consistency function $f_{\boldsymbol{\phi}}$ to obtain the one-step target sample $\boldsymbol{\theta}_{\varepsilon}=f_{\boldsymbol{\phi}}(\boldsymbol{\theta}_{T},T;\,\mathbf{x})$. What is more, inference with consistency model is not actually limited to one-step sampling. In fact, multi-step generation is possible with an iterative sampling procedure, which we will describe in the following. For a sequence of time points $\varepsilon=t_{1}<t_{2}<\dots<t_{K}=T$ and initial noise $\boldsymbol{\theta}_{K}\sim\mathcal{N}(\mathbf{0},T^{2}\mathbf{I})$, we calculate
@@
for $k=K-1,K-2,\dots,1$, where $\mathbf{z}_{k}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and $K-1$ is the number of sampling steps (Song & Dhariwal, 2023; Song et al., 2023). The resulting sample is usually better than a one-step sample: Adjusting the number of steps trades compute against sample quality (Song et al., 2023).
3.2 Consistency Models for SBI
Originally developed for image generation, consistency models can be applied to learn arbitrary distributions. The free-form architecture enables the integration of specialized architectures for both the data $\mathbf{x}$ and the parameters $\boldsymbol{\theta}$. Due to the low number of passes required for sampling (in contrast to flow matching and diffusion models), more complex networks can be used while maintaining low inference time. In theory, consistency models combine the best of both worlds (see Table 1): Unconstrained networks for optimal adaptation to parameter structure and data modalities, while enabling fast inference speed with few network passes. This comes at the cost of explicit invertibility, which leads to hardships for computing the posterior density, as detailed in Section 3.4. In accordance with the taxonomy from Cranmer et al. (2020), we call our method consistency model posterior estimation (CMPE). While we focus on posterior estimation, using consistency models for likelihood emulation is a natural extension of our work.
As a consequence of its fundamentally different training objective, CMPE is not just a faster version of FMPE. Instead, it shows qualitative differences to FMPE beyond a much faster sampling speed: (i) In Experiment 4, we show that CMPE is less dependent on the neural network architecture than FMPE, making it a promising alternative when the optimal architecture for a task is not known; (ii) throughout our experiments, we observe good performance in the low-data regime, rendering CMPE an attractive method when training data is scarce. In fact, limited data availability is a common limiting factor for complex simulation programs in science (e.g., molecular dynamics; Kadupitiya et al., 2020) and engineering (Heringhaus et al., 2022).
@@ 3.3 Optimization Objective
We use the consistency training objective for CMPE,
@@
where $\lambda(t)$ is a weighting function, $d(\mathbf{u},\mathbf{v})$ is a distance metric, $\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, and the arguments for the distance metric amount to the outputs of the consistency function obtained at two neighboring time indices $t_{i}$ and $t_{i+1}$,
$\mathbf{u}(\boldsymbol{\phi},t_{i+1};\mathbf{x})$ $=f_{\boldsymbol{\phi}}(\boldsymbol{\theta}+t_{i+1}\mathbf{z},t_{i+1};\,\mathbf{x})$ (8)
$\mathbf{u}(\boldsymbol{\phi}^{-},t_{i};\mathbf{x})$ $=f_{\boldsymbol{\phi}^{-}}(\boldsymbol{\theta}+t_{i}\mathbf{z},t_{i};\,\mathbf{x}).$ (9)
The teacher’s parameter vector $\boldsymbol{\phi}^{-}$ is a copy of the student’s parameter which is held constant during each step via a stopgrad operator, $\boldsymbol{\phi}^{-}\leftarrow\texttt{stopgrad}(\boldsymbol{\phi})$. We follow Song & Dhariwal (2023) using $\lambda(t_{i})=1/(t_{i+1}-t_{i})$ and $d(\mathbf{u},\mathbf{v})=\sqrt{\lVert\mathbf{u}-\mathbf{v}\rVert_{2}^{2}+c^{2}}-c$ from the Pseudo-Huber metric family. Appendix A contains details on discretization and noise schedules.
@@ 3.4 Density Estimation
where $f^{-1}_{\boldsymbol{\phi}}(\boldsymbol{\theta},T;\mathbf{x})$ is the implicit inverse of the consistency function and $\partial\boldsymbol{\theta}_{T}/\partial\boldsymbol{\theta}_{\varepsilon}$ is the resulting Jacobian. While we cannot explicitly evaluate $f^{-1}_{\boldsymbol{\phi}}(\boldsymbol{\theta}_{\varepsilon},T;\,\mathbf{x})$, we can generate samples from $\boldsymbol{\theta}_{T}\sim\mathcal{N}(\mathbf{0},T^{2}\mathbf{I})$ and use autodiff for evaluating the Jacobian because the consistency surrogate $f_{\boldsymbol{\phi}}$ is differentiable. Thus, we cannot directly evaluate the posterior density at arbitrary $\boldsymbol{\theta}$. Yet, evaluating the density at a set of $S$ posterior draws $\{\boldsymbol{\theta}_{\varepsilon}\}_{s=1}^{S}$ suffices for crucial downstream tasks, such as marginal likelihood estimation (Radev et al., 2023a), neural importance sampling (Dax et al., 2023a), or self-consistency losses (Schmitt et al., 2023a).
For the purpose of multi-step sampling, the latent variables $\boldsymbol{\theta}_{K},\boldsymbol{\theta}_{K-1},\dots,\boldsymbol{\theta}_{1}$ form a Markov chain with a change-of-variables (CoV) given by Köthe (2023):
@@ 3.5 Choosing the Number of Sampling Steps
The design of consistency models enables one-step sampling (see Eq. 6). In practice, however, using two steps significantly increases the sample quality in image generation tasks (Song & Dhariwal, 2023). We observe that few-step sampling with approximately $K=5-10$ steps provides the best trade-off between sample quality and compute, particularly in low-dimensional problems. This is roughly comparable to the speed of one-step estimators like affine coupling flows or neural spline flows in Experiments 1–3 (see Figure 3). The number of steps can be chosen at inference time, so practitioners can easily adjust this trade-off for a given situation. We noticed that approaching the maximum number of discretization steps used in training can lead to overconfident posterior distributions. This effect becomes even more pronounced when the maximum number of discretization steps during training is surpassed, which we therefore discourage based on our empirical results.
@@ 4 Empirical Evaluation
Our experiments cover three fundamental aspects of SBI. First, we perform an extensive evaluation on three low-dimensional experiments with bimodal posterior distributions from benchmarking suites for inverse problems (Lueckmann et al., 2021; Kruse et al., 2021). The simulation-based training phase is based on a fixed training set $\{(\mathbf{x}^{(m)},\boldsymbol{\theta}_{*}^{(m)})\}_{m=1}^{M}$ of $M$ tuples of data sets $\mathbf{x}^{(m)}$ and corresponding data-generating parameters (aka. ground-truths) $\boldsymbol{\theta}^{(m)}_{*}$. Second, we focus on an image denoising example which serves as a sufficiently high-dimensional case study in the context of SBI (Ramesh et al., 2022; Pacchiardi & Dutta, 2022). Third, we apply our CMPE method to a computationally challenging scientific model of tumor spheroid growth and showcase its superior performance for a non-trivial scientific simulator (Jagiella et al., 2017). We implement all experiments using the BayesFlow Python library for amortized Bayesian workflows (Radev et al., 2023b).
Evaluation metrics. We evaluate the experiments based on well-established metrics to gauge the accuracy and calibration of the results. All metrics are computed on a test set of $J$ unseen instances $\{(\mathbf{x}^{(j)},\boldsymbol{\theta}^{(j)}_{*})\}_{j=1}^{J}$. In the following, $S$ denotes the number of (approximate) posterior samples that we draw for each instance $J$. First, the easy-to-understand root mean squared error (RMSE) metric quantifies both the bias and variance of the approximate posterior samples across the test set as $\frac{1}{J}\sum_{j=1}^{J}\sqrt{\frac{1}{S}\sum_{s=1}^{S}\big{(}\boldsymbol{\theta}^{(j)}_{s}-\boldsymbol{\theta}_{*}^{(j)}\big{)}^{2}}$. Second, we estimate the squared maximum mean discrepancy (MMD; Gretton et al., 2012) between samples from the approximate vs. reference posterior, which is a kernel-based distance between distributions. Third, as a widely applied metric in SBI, the C2ST score uses an MLP classifier to distinguish samples from the approximate and reference posteriors. The resulting test accuracy of the classifier is the C2ST score, which ranges from 0.5 (worst; samples are indistinguishable) to 1 (best; samples are perfectly separable). Finally, uncertainty calibration is assessed through simulation-based calibration (SBC; (Talts et al., 2018)): All uncertainty intervals $U_{q}(\boldsymbol{\theta}\mid\mathbf{x})$ of the true posterior $p(\boldsymbol{\theta}\mid\mathbf{x})$ are well calibrated for every quantile $q\in(0,1)$,
@@
Contender methods. We compare affine coupling flows (ACF; Dinh et al., 2016), neural spline flows (NSF; Durkan et al., 2019), flow matching posterior estimation (FMPE; Dax et al., 2023b), and consistency model posterior estimation (CMPE; ours). We use identical free-form neural networks for FMPE and CMPE to ensure comparability.
Figure 3: Experiment 1 (Gaussian Mixture Model). Average C2ST and its standard deviation over a test set of size $J=100$, as a function of each method’s sampling speed for 1000 posterior draws. The sampling speed of FMPE and CMPE (Ours) is controlled by setting the number of inference steps $K$. CMPE (Ours) outperforms all other methods through both faster and better inference, as indexed by lower (better) C2ST scores. Mind the broken $x$-axis.
@@ 4.1 Experiment 1: Gaussian Mixture Model
We illustrate CMPE on a 2-dimensional Gaussian mixture model with two symmetrical components (Geffner et al., 2022; Schmitt et al., 2023a). The symmetrical components are equally weighted and have equal variance,
@@
where $\mathcal{N}(\cdot\,|\,\mathbf{\mu},\mathbf{\Sigma})$ is a Gaussian distribution with location $\mathbf{\mu}$ and covariance matrix $\mathbf{\Sigma}$. The resulting posterior distribution is bimodal and point-symmetric around the origin (see Figure 2 top-left). Each simulated data set $\mathbf{x}$ consists of ten exchangeable observations $\{\mathbf{x}_{1},\ldots,\mathbf{x}_{10}\}$, and we train all architectures on $M=1024$ training simulations. We use a DeepSet (Zaheer et al., 2017) to learn 6 summary statistics for each data set jointly with the inference task.
Results. As displayed in Figure 2, both ACF and NSF fail to fit the bimodal posterior with separated modes. FMPE successfully forms disconnected posterior modes, but shows visible overconfidence through overly narrow posteriors. The visual sampling performance of CMPE is superior to all other methods. In this task, we observe that CMPE does not force us to choose between speed or performance relative to the other approximators. Instead, CMPE can outperform all other methods simultaneously with respect to both speed and performance, as evidenced by lower C2ST to the reference posterior across $J=100$ test instances (see Figure 3). If we tolerate slower sampling, CMPE achieves peak performance at $K=10$ inference steps. Most notably, CMPE outperforms 1000-step FMPE by a large margin, even though the latter is approximately 75$\times$ slower.
Figure 4: Experiment 2 (Two Moons). CMPE (Ours) with 10 sampling steps shows superior performance up to a training budget of 4096 instances. Plots show average C2ST and its standard error over a test set of size $J=100$.
Figure 5: Experiment 4. CMPE denoising results from each class of Fashion MNIST obtained using a U-Net architecture and two-step sampling. A training set of 60 000 images was used. First row: Original image acting as the “parameters” of the noisy camera simulator. Second row: Blurred image, acting as the output of the camera simulator. Third and fourth row: Means and standard deviations of the posteriors estimated from the corresponding blurry “observations”. Note: For standard deviations, darker regions indicate larger variability in the outputs. Adapted from Radev et al. (2023a).

### v3

@@ 3 Consistency model posterior estimation
Diffusion models have one crucial drawback: At inference time, they require solving a differential equation for each posterior draw which slows down their sampling speed. This is particularly troublesome in SBI applications which may require thousands of samples for thousands of data sets [e.g., $>$1M data sets in 7]. Consistency models [CMs; 18] address this problem with a new generative model family that supports both single-step and multi-step sampling. In the following, we summarize key ideas of CMs [18, 46] with a focus on conditional sampling for the purpose of SBI.
@@ 3.1 Conditional consistency models
The consistency function $f:(\boldsymbol{\theta}_{t},t;\,\mathbf{x})\mapsto\boldsymbol{\theta}_{\varepsilon}$ maps points across the solution trajectory $\{\boldsymbol{\theta}_{t}\}_{t\in[T,\varepsilon]}$ to the trajectory’s origin $\boldsymbol{\theta}_{\varepsilon}$ given a fixed conditioning variable (i.e., observation) $\mathbf{x}$ and the probability flow ODE in Eq. 5. To achieve this with established score-based diffusion model architectures, we can use a unconstrained neural network $F_{\boldsymbol{\phi}}(\boldsymbol{\theta},t;\,\mathbf{x})$ which is parameterized through skip connections of the form
@@
where $c_{\text{skip}}(t)$ and $c_{\text{out}}(t)$ are differentiable and fulfill the boundary conditions $c_{\text{skip}}(\varepsilon)=1$ and $c_{\text{out}}(\varepsilon)=0$. Consistency models are originally motivated as a distillation technique for diffusion models. However, Song et al. [18] show that training consistency models in isolation is possible, and we base our method on their direct approach.
Sampling. Once the consistency model has been trained, generating draws from the approximate posterior is straightforward by drawing samples from the noise distribution, $\boldsymbol{\theta}_{T}\sim\mathcal{N}(\mathbf{0},T^{2}\mathbf{I})$, which shall then be transformed into samples from the target distribution, like in a standard diffusion model. In contrast to diffusion models, however, we do not need to solve a sequence of differential equations for this transformation. Instead, we can use the learned consistency function $f_{\boldsymbol{\phi}}$ to obtain the one-step target sample $\boldsymbol{\theta}_{\varepsilon}=f_{\boldsymbol{\phi}}(\boldsymbol{\theta}_{T},T;\,\mathbf{x})$. What is more, inference with consistency models is not actually limited to one-step sampling. In fact, multi-step generation is possible with an iterative sampling procedure, which we will describe in the following. For a sequence of time points $\varepsilon=t_{1}<t_{2}<\dots<t_{K}=T$ and initial noise $\boldsymbol{\theta}_{K}\sim\mathcal{N}(\mathbf{0},T^{2}\mathbf{I})$, we calculate
@@
for $k=K-1,K-2,\dots,1$, where $\mathbf{z}_{k}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and $K-1$ is the number of sampling steps [46, 18]. The resulting sample is usually better than a one-step sample.
3.2 Consistency models for simulation-based inference
Originally developed for image generation, consistency models can be applied to learn arbitrary distributions. The unconstrained architecture enables the integration of specialized architectures for both the data $\mathbf{x}$ and the parameters $\boldsymbol{\theta}$. Due to the low number of passes required for sampling (in contrast to flow matching and diffusion models), more complex networks can be used while maintaining low inference time. In theory, consistency models combine the best of both worlds: unconstrained networks for optimal adaptation to parameter structure and data modalities, while enabling fast inference speed with few network passes. Currently, this comes at the cost of explicit invertibility, which limits the computation of posterior densities. More precisely, single-step consistency models do not allow density evaluations at an arbitrary parameter value $\boldsymbol{\theta}$ but only at a set $S$ of approximate posterior draws $\{\boldsymbol{\theta}_{\varepsilon}^{(1)},\dots,\boldsymbol{\theta}_{\varepsilon}^{(S)}\}$. However, this is sufficient for important downstream tasks like marginal likelihood estimation, importance sampling, or self-consistency losses. In contrast, multi-step consistency sampling defines a Markov chain which cannot be evaluated without an additional density estimator (see Section 3.5 for details). In accordance with the taxonomy from Cranmer et al. [1], we call our method consistency model posterior estimation (CMPE). While we focus on posterior estimation, using consistency models for likelihood emulation is a natural extension of our work.
As a consequence of its fundamentally different training objective, CMPE is not just a faster version of FMPE. Instead, it shows qualitative differences to FMPE beyond a much faster sampling speed. In Experiment 4, we show that CMPE is less dependent on the neural network architecture than FMPE, making it a promising alternative when the optimal architecture for a task is not known. Further, we consistently observe good performance in the low-data regime, rendering CMPE an attractive method when training data is scarce. In fact, data availability is a common limiting factor for complex simulators in engineering [47] and science [e.g., molecular dynamics; 48].
@@ 3.3 Optimization objective
We formulate the consistency training objective for CMPE, which extends the unconditional training objective from Song and Dhariwal [46] with a conditioning variable $\mathbf{x}$ to cater to the SBI setting,
@@
where $\lambda(t)$ is a weighting function, $d(\mathbf{u},\mathbf{v})$ is a distance metric, $\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is unit Gaussian noise, and the arguments for the distance metric amount to the outputs of the consistency function obtained at two neighboring time indices $t_{i}$ and $t_{i+1}$,
$\mathbf{u}(\boldsymbol{\phi},t_{i+1};\mathbf{x})=f_{\boldsymbol{\phi}}(\boldsymbol{\theta}+t_{i+1}\mathbf{z},t_{i+1};\,\mathbf{x}),\;\;\mathbf{u}(\boldsymbol{\phi}^{-},t_{i};\mathbf{x})=f_{\boldsymbol{\phi}^{-}}(\boldsymbol{\theta}+t_{i}\mathbf{z},t_{i};\,\mathbf{x}).$ (9)
The teacher’s neural network weights $\boldsymbol{\phi}^{-}$ are a copy of the student’s weights which are held constant during each step via a stopgrad operator, $\boldsymbol{\phi}^{-}\leftarrow\texttt{stopgrad}(\boldsymbol{\phi})$. We follow Song and Dhariwal [46] using $\lambda(t_{i})=1/(t_{i+1}-t_{i})$ and $d(\mathbf{u},\mathbf{v})=\sqrt{\lVert\mathbf{u}-\mathbf{v}\rVert_{2}^{2}+c^{2}}-c$ from the Pseudo-Huber metric family. Appendix A contains more details on discretization and noise schedules.
3.4 Hyperparameter tuning
Consistency training introduces several additional hyperparameters, but there is limited theoretical guidance on how to select most of them. Within the scope of this paper, we rely on empirical search to identify the hyperparameters that are the most relevant for tuning efforts. As stated above, the maximum time $T$ determines the standard deviation of the latent distribution. It should be larger than the magnitude of the target distribution [18]. Values that are too low might lead to biased sampling, while values that are too high can make the training process more difficult. The minimum and maximum number of discretization steps $s_{0},s_{1}$ during training also influence the final result in our experiments. For $s_{0}$, a sufficiently low value (e.g. $10$) should be chosen. For $s_{1}$, smaller values around $50$ seem beneficial for stable training, especially when the number of epochs is low. When longer training is possible, increasing $s_{1}$ might lead to improved accuracy. We summarize our recommended hyperparameter choices for SBI applications in Appendix C.
@@ 3.5 Density estimation
where $f^{-1}_{\boldsymbol{\phi}}(\boldsymbol{\theta},T;\mathbf{x})$ is the implicit inverse of the consistency function and $\partial\boldsymbol{\theta}_{T}/\partial\boldsymbol{\theta}_{\varepsilon}$ is the resulting Jacobian. While we cannot explicitly evaluate $f^{-1}_{\boldsymbol{\phi}}(\boldsymbol{\theta}_{\varepsilon},T;\,\mathbf{x})$, we can generate samples from $\boldsymbol{\theta}_{T}\sim\mathcal{N}(\mathbf{0},T^{2}\mathbf{I})$ and use autodiff for evaluating the Jacobian because the consistency surrogate $f_{\boldsymbol{\phi}}$ is differentiable. Thus, we cannot directly evaluate the posterior density at arbitrary $\boldsymbol{\theta}$. Yet, evaluating the density at a set of $S$ posterior draws $\{\boldsymbol{\theta}_{\varepsilon}\}_{s=1}^{S}$ suffices for crucial downstream tasks, such as marginal likelihood estimation [49], neural importance sampling [21], or self-consistency losses [50].
For the purpose of multi-step sampling, the latent variables $\boldsymbol{\theta}_{K},\boldsymbol{\theta}_{K-1},\dots,\boldsymbol{\theta}_{1}$ form a Markov chain with a change-of-variables (CoV) given by Köthe [51]:
@@ 3.6 Choosing the number of sampling steps
The design of consistency models enables one-step sampling (see Eq. 7). In practice, however, using two steps significantly increases the quality in image generation tasks [46]. We observe that few-step sampling with approximately $K=5{-}15$ steps provides the best trade-off between sample quality and compute for SBI, particularly in low-dimensional problems. This is roughly comparable to the speed of neural one-step estimators like affine coupling flows or neural spline flows in Experiments 1–3 (see 2(a)). The number of steps can be chosen at inference time, so practitioners can easily adjust this for a given situation. We noticed that approaching the maximum trained number of discretization steps can lead to overconfident posterior distributions. See Section C.5 for an empirical evaluation of the relation between sampling quality and number of sampling steps.
@@ 4 Empirical evaluation
Our experiments cover three fundamental aspects of SBI. First, we perform an extensive evaluation on three low-dimensional experiments with bimodal posterior distributions from benchmarking suites for inverse problems [52, 53]. The simulation-based training phase is based on a fixed training set $\{(\mathbf{x}^{(m)},\boldsymbol{\theta}_{*}^{(m)})\}_{m=1}^{M}$ of $M$ tuples of data sets $\mathbf{x}^{(m)}$ and corresponding ground-truth parameters $\boldsymbol{\theta}^{(m)}_{*}$. Second, we focus on an image denoising example which serves as a sufficiently high-dimensional case study in the context of SBI [54, 32]. Third, we apply our CMPE method to a computationally challenging scientific model of tumor spheroid growth and showcase its superior performance for a complex simulator from cell biology [19]. We implement all experiments using the BayesFlow Python library for amortized Bayesian workflows [55].
Evaluation metrics. We evaluate the experiments based on well-established metrics to gauge the accuracy and calibration of the results. All metrics are computed on a test set of $J$ unseen instances $\{(\mathbf{x}^{(j)},\boldsymbol{\theta}^{(j)}_{*})\}_{j=1}^{J}$. In the following, $S$ denotes the number of (approximate) posterior samples that we draw for each instance $J$. First, the easy-to-understand root mean squared error (RMSE) metric quantifies both the bias and variance of the approximate posterior samples across the test set as $\frac{1}{J}\sum_{j=1}^{J}\sqrt{\frac{1}{S}\sum_{s=1}^{S}\big{(}\boldsymbol{\theta}^{(j)}_{s}-\boldsymbol{\theta}_{*}^{(j)}\big{)}^{2}}$. Second, we estimate the squared maximum mean discrepancy [MMD; 56] between samples from the approximate vs. reference posterior, which is a kernel-based distance between distributions. Third, as a widely applied metric in SBI, the C2ST score uses an MLP classifier to distinguish samples from the approximate and reference posteriors. The resulting test accuracy of the classifier is the C2ST score, which ranges from 1 (samples are perfectly separable $\rightarrow$ bad posterior) to 0.5 (samples are indistinguishable $\rightarrow$ perfect posterior). Finally, we assess uncertainty calibration through simulation-based calibration (SBC; [57]): All uncertainty intervals $U_{q}(\boldsymbol{\theta}\mid\mathbf{x})$ of the true posterior $p(\boldsymbol{\theta}\mid\mathbf{x})$ are well calibrated for every quantile $q\in(0,1)$,
@@
Contender methods. We compare affine coupling flows [ACF; 40], neural spline flows [NSF; 41], flow matching posterior estimation [FMPE; 11], and consistency model posterior estimation (CMPE; ours). We use identical unconstrained neural networks for FMPE and CMPE to ensure comparability. Appendix C lists the hyperparameters of all models in the experiments.
(a) Experiment 1 (Gaussian Mixture Model)
(b) Experiment 2 (Two Moons)
Figure 2: C2ST score of 4000 approximate posterior draws vs. reference posterior (lower is better) for $J=100$ unseen test examples. (2(a)) CMPE (Ours) outperforms all other methods through both faster and more accurate inference on the GMM benchmark (mean±$\pm$±SD). (2(b)) CMPE (Ours) with 10 sampling steps shows superior performance up to a training budget of 4096 instances on the Two Moons benchmark (mean±$\pm$±SE).
@@ 4.1 Experiment 1: Gaussian mixture model
We illustrate CMPE on a 2-dimensional Gaussian mixture model with two symmetrical components [9, 50]. The symmetrical components are equally weighted and have equal variance,
@@
where $\mathcal{N}(\cdot\,|\,\mathbf{\mu},\mathbf{\Sigma})$ is a Gaussian distribution with location $\mathbf{\mu}$ and covariance matrix $\mathbf{\Sigma}$. The resulting posterior distribution is bimodal and point-symmetric around the origin (see Figure 1 top-left). Each simulated data set $\mathbf{x}$ consists of ten exchangeable observations $\{\mathbf{x}_{1},\ldots,\mathbf{x}_{10}\}$, and we train all architectures on $M=1024$ training simulations. We use a DeepSet [58] to learn 6 summary statistics for each data set jointly with the inference task.
Results. As displayed in Figure 1, both ACF and NSF fail to fit the bimodal posterior with separated modes. FMPE successfully forms disconnected posterior modes, but shows visible overconfidence through overly narrow posteriors. The visual sampling performance of CMPE is superior to all other methods. In this task, we observe that CMPE does not force us to choose between speed or performance relative to the other approximators. Instead, CMPE can outperform all other methods simultaneously with respect to both speed and performance, as quantified by lower C2ST to the reference posterior across $J=100$ test instances (see 2(a)). If we tolerate slower sampling, CMPE achieves peak performance at $K=10$ inference steps. Most notably, CMPE outperforms 1000-step FMPE by a large margin, even though the latter is approximately 75$\times$ slower.